A rocket is launched with velocity 10 km/s. If the radius of earth is R, then maximum height attained by it will be 
(a) 2R        (b)3R     (ć)4R    (d) 5R

Grade:12th pass

1 Answers

Sai
5 years ago
KE of body=PE of body
1/2mv^2=mg(R+H)
Where H is maximum height
V^2=2g(R+H)
(10*5000/8)^2=20(R+H)
SINCE R=6.4*10^6
5R=R+H
H=4R
SOLVING WE GET